For the following reaction, 4.41 grams of hydrogen gas are allowed to react with 39.8 grams of iodine .


hydrogen(g) + iodine(s)  hydrogen iodide(g)


What is the maximum mass of hydrogen iodide that can be formed?  grams


What is the FORMULA for the limiting reagent?


What mass of the excess reagent remains after the reaction is complete?  grams


1
Expert's answer
2022-05-26T15:17:03-0400

H2 + I2 = 2HI

n(H2) = m(H2)/Mr(H2) = 4.41 g / 2 g/mol = 2.2 moles

n(I2) = m(I2) / Mr(I2) = 39.8 g / 254 g/mol = 0.16 moles

H2 - excess reagent

I2 - limiting reagent


m(HI) = (m(I2) * 2Mr(HI)) / Mr(I2) = (39.8 g * 2*128 g/mol) / 254 g/mol = 40.11 g


m(H2) = (m(I2)*Mr(H2)) / Mr(I2) = (39.8 g * 2 g/mol) / 254 g/mol = 0.31 g

after a the reaction m(H2) = 4.41 g - 0.31 g = 4.1 g